OpenStudy (anonymous):

An observer (O) spots a plane flying at a 42° angle to his horizontal line of sight. If the plane is flying at an altitude of 15,000 ft., what is the distance (x) from the plane (P) to the observer (O)? A right triangle is shown with one angle marked 42 degrees, one side marked x and the height marked 15000 feet. 10,035 feet 16,648 feet 20,188 feet 22,417 feet

OpenStudy (jdoe0001):

got picture?

OpenStudy (anonymous):

idk how tp post it

OpenStudy (jdoe0001):

use the blue button of [Attach File] below this text

OpenStudy (anonymous):

ok got it ;)

OpenStudy (jdoe0001):

 \(\bf sin(\theta)=\cfrac{opposite}{hypotenuse}\implies hypotenuse=\cfrac{opposite}{sin(\theta)}\)

OpenStudy (anonymous):

so would i put 1500o over x because that is the hypotenuse

OpenStudy (jdoe0001):

well..yes... but you would want to find "x".... so \(\bf sin(\theta)=\cfrac{\frac{opposite}{(15,000)}}{\frac{hypotenuse}{x}}\implies \frac{hypotenuse}{x}=\cfrac{\frac{opposite}{(15,000)}}{sin(\theta)}\)

OpenStudy (jdoe0001):

recall, your hypotenuse is "x" and your angle is 42 degrees and your opposite side is 15,000

OpenStudy (jdoe0001):

\(\bf sin(\theta)=\cfrac{opposite}{hypotenuse}\to hypotenuse=\cfrac{opposite}{sin(\theta)}\to x=\cfrac{15,000}{sin(42^o)}\)

OpenStudy (anonymous):

ok im starting to understand but idk how to type this into the calculator i know the functions i just dont know how to place it

OpenStudy (jdoe0001):

42 and then press [sin] button make sure your calculator is in Degree mode

OpenStudy (anonymous):

I think that plane is 16659.188 feet away - see picture: http://www.triangle-calculator.com/?what=rt&a=A%3D42+a%3D15000&submit=Solve

OpenStudy (jdoe0001):

how about using -> https://desmos.com/calculator

OpenStudy (anonymous):

ok i did what u said but the numbers cant be rite ill try desmos hold on and thank you for taking time to help me

OpenStudy (anonymous):

ok i got it ill try one omo and can u check it for me pls?

OpenStudy (jdoe0001):

well. you can always post anew... in case I dunno, someone else may and we can revise each other :)

OpenStudy (anonymous):

ok thanks ; )

OpenStudy (anonymous):

A bird (B) is spotted flying 6,000 feet from a tower (T)). An observer (O) spots the top of the tower (T)) at a distance of 9,000 feet. What is the angle of depression from the bird (B) to the observer (O)? Right triangle OTB is shown. Side OT labeled 9000 and side TB is labeled 9000. The angle B is labeled x degrees. 33.69° 41.81° 48.18° 56.31°

OpenStudy (anonymous):

would i use tangent